Mathematical Models with Boundary Value

Problems

Douglas B. Meade  Allan A. Struthers
Department of Mathematics  Department of Mathematical Sciences
University of South Carolina  Michigan Technical University
Columbia, SC 29208  Houghton, MI 49931
(803) 777-6183  (906) 487-3541

     meade@math.sc.edu                 struther@math.mtu.edu
http://www.math.sc.edu/˜ meade/    http://www.math.mtu.edu/˜ struther/

8 November 1997